Monday, July 18, 2011

Week 8

Last week was a busy week, which kept me from posting much. I'll try to give a somewhat decent account of what happened, and the status of our projects. Last week, we had some trouble getting consistent results from Fash's buffers. We investigated a variety of variables, from my calibration technique to pippetting errors. We also had a problem with counts from the spectrometer. We eventually discovered that part of our problem may have been coming from the plasma torch area being dirty. There was also a possible problem with consistent nitrogen flow. We now think we have most of the issues figured out, and should be able to proceed with collecting consistent results. Ben has been getting consistent buffer values, and is proceeding with his NCP experiments. For my project, last week I ran the second buffer solution, and began to test the actual samples. The second buffer solution had somewhat surprising concentrations, because they came out to consistently around 8 mM Cohex, when we thought they were 10. I am going to double check these results after I test the actual samples further.

No comments:

Post a Comment